# Concurrent-Learning Based Relative Localization in Shape Formation of Robot Swarms

### v1

@@ Abstract
In this paper, we address the shape formation problem for massive robot swarms in environments where external localization systems are unavailable. Achieving this task effectively with solely onboard measurements is still scarcely explored and faces some practical challenges. To solve this challenging problem, we propose the following novel results. Firstly, to estimate the relative positions among neighboring robots, a concurrent-learning based estimator is proposed. It relaxes the persistent excitation condition required in the classical ones such as least-square estimator. Secondly, we introduce a finite-time agreement protocol to determine the shape location. This is achieved by estimating the relative position between each robot and a randomly assigned seed robot. The initial position of the seed one marks the shape location. Thirdly, based on the theoretical results of the relative localization, a novel behavior-based control strategy is devised. This strategy not only enables adaptive shape formation of large group of robots but also enhances the observability of inter-robot relative localization. Numerical simulation results are provided to verify the performance of our proposed strategy compared to the state-of-the-art ones. Additionally, outdoor experiments on real robots further demonstrate the practical effectiveness and robustness of our methods.
@@
Shape formation has a broad potential for large group of robots to execute certain tasks, such as object transport, forest firefighting, and entertainment shows. However, most of the existing approaches rely on external localization infrastructures, rendering them impractical in environments where such systems are not available. To address this issue, this paper proposes an integrated strategy that can achieve shape formation for large group of robots by using local distance and displacement measurements. This strategy consists of three main components. Firstly, a relative localization estimator is introduced to estimate the relative positions among neighboring robots. Secondly, a protocol for reaching a consensus on the desired shape’s position is proposed. Thirdly, a behavior-based controller is developed to achieve massive shape formation and enhance the observability of relative localization. More details of the proposed algorithms and swarm robotic systems are provided in this paper.
@@ I Introduction
Shape formation has received significant attention in recent years due to its broad potential across many applications [1, 2, 3]. The objective of shape formation is to guide the robot swarms from an initial configuration to form a user-specified shape through local interactions. To achieve this objective, various control methods have been explored, including the ones based on artificial potential [4, 5], edge following [6, 7], mean-shift exploration [8, 9], and goal assignment [10, 11]. However, most of the methods rely on external localization infrastructures like Global Position System (GPS) or Beidou Navigation Satellite System (BDS), to localize robots, which can be restrictive for practical applications [12]. For instance, external localization systems may be impractical or infeasible in some scenarios, such as indoor environment, underwater, or building atrium. Rather than relying on external localization infrastructures, robots equipped solely with onboard sensors are more autonomous.
@@
In this paper, we aim to investigate the problem of relative localization in shape formation using only onboard measurements, as shown in Fig. 1. To handle this problem, the existing results can be roughly classified into the following two categories. The first class is based on the rigidity conditions. These methods usually require an assumption that the measurement topology among robots satisfy rigidity conditions like distance rigidity [13, 14] or bearing rigidity [15, 16]. The second class, which has also attracted research attention, is rigidity-free methods. To that end, researchers incorporate the persistent excitation (PE) characteristics into the control command, such as noise [17, 18] or time-varying reference [19, 12], to relax rigidity conditions on the measurement topology. Although rigidity and PE conditions are necessary in some collaborative tasks, they may not be essential in the context of massive shape formation. More significantly, imposing these conditions may bring a series of technical challenges, as outlined below.
@@
Firstly, due to the rigidity condition, the two robots must maintain continuous wireless communication and onboard measurements if they are topological neighbors in the desired formation. Additionally, it is necessary for the measurement topology of the entire robot swarms to satisfy the rigidity condition. This requirement is the foundation of the global convergence property of most existing results [13, 14, 15, 16]. However, the two robots that are topological neighbors may be far apart physically, which bring challenges for measurement and communication between them. Furthermore, the rigidity condition across the entire topology is also difficult to guarantee, especially for large group of robots.
@@
Secondly, to relax the rigidity condition, researchers introduce the PE condition based localization estimator. But existing methods often involve sacrifices or encounter new limitations in order to satisfy this condition. In practice, the most common way to fulfill the PE condition is embedding noise or time-varying reference into the control commands of individual robots [17, 18]. However, these PE characteristics may propagate within robot swarms due to interactions between robots. This can result in non-smooth motion trajectories for robot swarms, leading to deteriorated control performance in real-world applications, particularly with large group of robots. Although the PE condition can be met through specific movements like rotation [19] and spinning [12], these methods are ineffective for common static shape formation control.
@@
Motivated by these challenges, we focus on the shape formation control for large group of robots without PE condition. Specifically, the contribution of this paper are summarized below.
@@
1) To achieve inter-robot relative localization without the PE condition, a concurrent-learning based relative position estimator is introduced. Compared with the existing results [18, 17, 19], both current and historical measurement data are utilized in the estimator. As a result, it is not necessary to incorporate PE characteristics into the control command, resulting in smoother motion trajectories for the robots. Theoretical analysis results are provided to guarantee the convergence of the estimation error. Furthermore, we provide the optimal strategy for relative motion and data collection between robots. We demonstrate that the proposed estimator achieves the largest convergence rate with the optimal strategy.
@@
2) To achieve the shape localization agreement for the entire swarms in the absense of global coordinates, the location of the shape center is selected as the initial position of a randomly assigned seed robot. Note that the position of the seed robot is irrelevant. By transforming the global-coordinate based shape agreement problem [8] into a relative position estimation problem, we propose a finite-time consensus-based agreement protocol. This approach eliminates the need for the rigidity condition on the measurement topology [13, 14, 15, 16] to achieve shape localization agreement. Furthermore, our method does not necessitate knowledge of the absolute position of the seed robot. Theoretical analysis results are also provided to guarantee the convergence of the proposed method.
@@
3) For the shape formation control of the robot swarms, we design a novel behavior-based control strategy which is based on the theoretical insights from relative localization. This strategy enables shape formation in large group of robots in the absense of external localization system. Additionally, it enhances the observability of relative localization among robots. Extensive simulation results are provided to verify the effectiveness of the proposed methods. Moreover, compared with the existing swarm shape formation methods based on relative localization [12, 20, 21], outdoor experiments with number of robot up to six are also provided to verify the effectiveness, stability, and robustness of the proposed method.
@@
The rest of this paper is organized as follows. Section II provides the related works. Section III introduces an overview of the entire system, adaptive relative localization method and behavior-based shape formation control strategy are proposed in Section IV and Section V. Simulation and experimental results are discussed in Section VI and Section VII. Finally, Section VIII concludes this paper.
@@ II-A Distributed Shape Formation
Traditional researches on swarm formation control rely on consensus and graph theory [22, 23, 24]. In these approaches, each robot achieves shape formation by tracking a predefined desired reference, either in a centralized or distributed manner. However, the neighboring robots defined in the desired formation may be far apart physically, which is unreasonable in real applications.
Inspired by the phenomenon in nature where groups of insects and animals can form various spatial shapes through local interactions among individuals [25, 26, 27], extensive behavior-based studies on shape formation of robot swarms are proposed. The existing behavior-based control methods for shape formation can be roughly divided into two classes. One class is based on goal assignment. In these methods, each robot is dynamically assigned a unique goal location using either centralized approaches such as network-flow algorithms [28] or decentralized approaches such as distributed auction [10] and task-swapping algorithm [11]. Once the robots are assigned unique goal locations corresponding to the desired shape, the subsequent task is simply to guide the robots to reach their respective goal locations. Many existing results such as distributed model predictive control [10], optimal reciprocal collision avoidance [29] and planning based method [30] can accomplish this task effectively.
The other class of strategies are free of assignment. Inspired by natural processes like embryogenesis, the method in [31] can spontaneously generate swarm shapes with a reaction-diffusion network. Similarly, a method combining random walk and distributed edge detection is proposed in [2]. These methods exhibit significant randomness in both the shape formed and the robots’ motion trajectories. Another common assignment-free methods for shape formation are based on artificial potential fields (APF). However, APF-based methods often encounter a well-known locally minimum problem. To address this issue, an information integration and feedback based method is proposed in [5]. However, relying solely on APF cannot effectively avoid the deadlock phenomena. In our previous work [8], the idea of mean-shift exploration is adopted in shape formation of robot swarms, which significantly improves the efficiency of shape formation, especially for large group of robots.
In summary, all these methods are based on the global coordinates of robots. However, one can see that many simple organisms in nature can achieve complex shapes with only basic measurement information rather than global coordinates. Consequently, the development of shape formation techniques for robot swarms using only local measurements remains significant.
@@ II-B Relative Localization Estimation
Relative localization between robots is a fundamental problem in swarm robot research. Unlike traditional anchor-based localization methods [32, 33], we focus on the relative localization strategy which do not rely on external anchors. In recent years, visual-measurement based relative localization schemes have received significant attention. The work in [34] proposes a fisher information matrix (FIM) based localization-enhanced method with bearing measurement. To fully utilize visual measurement information and improve localization accuracy, a bearing-angle based localization strategy is proposed in [35]. However, the limited camera’s field of view may constrain the flexibility of the robot. Besides, the target detection algorithms used in bearing measurement may lead to high computational resource consumption.
In contrast, thanks to the availability of low-cost Ultra Wide Band (UWB) ranging sensors, distance based localization schemes are also widely explored. A common approach is the multiple anchors based localization strategy. In [36] and [37], multiple UWB tags are equipped on a robot, and relative localization is achieved using optimization methods like nonlinear least squares. However, the use of multiple UWB sensors may increase the structural complexity of the robot. In [18], a distance and displacement measurement based relative position estimator is proposed, requiring only one UWB sensor. However, this estimator only converges when the relative motion trajectory between robots satisfies the PE condition. As a result, periodic noise signal is added to the control command, which is similar to [17]. In [20], based on the FIM theory, the authors analyse how the relative motion between two robots can enhance relative localization. In the latest result [38], the PE condition is satisfied through the rotation of a UWB module on each robot. Nevertheless, the design of the turntable adds complexity to the robot’s structure.
@@ II-C Local Measurement Based Shape Formation
To achieve shape formation tasks with only local measurements, researchers in control and robotics have explored various strategies.
One common strategy is based on trilateral localization algorithm. In [6], a distance measurement based shape formation strategy is proposed. In this work, each robot localizes itself by measuring distance with at least four localized robots. However, this method requires the placement of anchor robots, increasing the difficulty of real-world application. Moreover, only edge robots can move to form shapes, resulting in lower efficiency. Similarly, multiple ranging sensors are used for inter-robot localization in [36]. In [39], elastic link is adopted to enhance localization for error-prone individuals. Despite their merits, these methods rely on external facilities, making practical implementation challenging.
@@
In summary, while the existing approaches offer promising solutions for local measurement based swarm shape formation control, they all have limitations. Currently, there is still no method that achieves efficient shape formation control for large group of robots without relying on the PE condition. The distributed shape formation problem with solely onboard measurements remains a challenge yet to be fully addressed.
III System Overview
III-A Notions and Problem Statement
@@
Consider a group of $n+1$ mobile robots in $\mathbb{R}^{2}$ and $n\geq 2$. Let $p_{i}\in\mathbb{R}^{2}$ be the position of robot $i=0,1,\ldots,n$ in an inertial frame. The dynamic model of each robot is considered as $\dot{p}_{i}=v_{i}$, where $v_{i}$ is the velocity command to be designed. The velocity command $v_{i}$ satisfies $\lVert v_{i}\rVert\leq v_{\max}$, where $v_{\max}$ is the maximum velocity of each robot. The interaction network among the robots is described by an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, which is composed of a vertex set $\mathcal{V}=\{0,1,\ldots,n\}$ and an edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. If $(i,j)\in\mathcal{E}$, the two robots can receive information with each other and robot $j$ is a neighbor of robot $i$. The neighbor set of robot $i$ is denoted as $\mathcal{N}_{i}=\{j\in\mathcal{V}:(i,j)\in\mathcal{E}\}$. The detailed definition of the neighbor set $\mathcal{N}_{i}$ can be found in V-B.
@@
The objective of each robot is to achieve shape formation for the entire robot swarms through local distance and displacement measurements. To achieve this, we design an integration strategy including relative position estimation, shape localization agreement, and shape formation control. The overall system architecture is illustrated in Fig. 2. The first step involves designing a concurrent-learning based estimator to estimate relative positions to neighboring robots (see Fig. 2A and Section IV).
The second step is to design an agreement protocol to locate the desired shape which is marked by the initial position of the seed robot (see Fig. 2B and Section V-B). The final step is to devise a behavior-based control law to achieve shape formation task and, meanwhile, enhance the localization (see Fig. 2C and Section V-C).
III-B Relative Localization to Neighbors
@@
To achieve independence from external localization infrastructure, robots first need to establish relative localization with respect to their neighbors using only onboard measurements. This paper enables each robot to estimate the relative position to its neighbors solely through inter-robot distance and displacement measurements (see Fig. 2A) which can be obtained using the low-cost sensors such as UWB and inertial odometry (IO), respectively.
The relative localization problem can be formulated as an adaptive parameter estimation problem [17]. Existing works for localization estimation often rely on embedding the PE characteristics to ensure the convergence of the parameter estimator [12, 17]. However, implementing PE typically involves adding noise or disturbance to the control input, compromising control performance [40]. In contrast, we propose distributed adaptive estimators for relative position estimation among neighboring robots based on the concept of concurrent-learning: robot leverages historical and present measurement data to alleviate the need for the PE condition (see Fig. 2A). In this way, there is no requirement to add persistent noise to the control command, resulting in smoother motion trajectories during shape formation. Theoretical and experimental results demonstrate that exponential convergence of the estimation error is achieved and the motion trajectories of robots are smoother than methods using persistent disturbance [18].
III-C Localization Agreement on Shape
Relying solely on relative localization is insufficient to determine the localization of the desired shape to be formed by robots. Therefore, resolving how robots can determine the shape position using only local information remains a challenge. To address this, we propose a localization agreement protocol that enables robots to negotiate distributively and reach a consensus on the shape position, as shown in Fig. 2B.
@@ III-C Localization Agreement on Shape
To initiate the shape formation, a robot is randomly designated as the seed robot, whose initial position marks the localization where the desired shape should be formed. Without losing generality, we denote the ID of the seed robot as 00. The seed robot plays a stubborn role by insisting on propagating the relative displacement from its initial position (i.e., shape position) to its non-seed neighbors. These non-seed robots utilize this propagated relative displacement to interpret the shape position locally through the proposed agreement protocol, simultaneously propagating their interpretations to other non-seed robots. In this way, the interpretations of non-seed robots gradually converge to the seed one so that all the robots reach a consensus on the localization of the desired shape with the proposed agreement protocol (see Fig. 2B and Section V-B. Theoretical results demonstrate that the proposed agreement protocol guarantees the finite-time convergence of shape localization among robots.
III-D Distributed Formation Shape Control
@@
Once all the robots reach an agreement on the localization of the desired shape, each robot can form the desired shape driven by the proposed behavior-based control law (see Section V). Specifically, the behavior-based control law consists of four velocity commands $v_{i}^{\rm ent}$, $v_{i}^{\rm exp}$, $v_{i}^{\rm int}$, and $v_{i}^{\rm enh}$. The four commands represent the shape-entering, shape-exploring, neighbor-interacting, and localization-enhancing behavior motions, respectively. The detailed usages of these command are as follows.
The shape-entering command $v_{i}^{\rm ent}$ guides robot $i$ into the desired shape by seeking the darkest cells around it (see Fig. 2C). The shape-exploring command $v_{i}^{\rm exp}$ aims to steer robot $i$ to explore the unoccupied area inside the desired shape (see Fig. 2C). The neighbor-interacting command $v_{i}^{\rm int}$ is responsible for both collision avoidance and velocity alignment among adjacent robots (see Fig. 2C). Different from our previous work [8] which relies on the external localization infrastructure, the above commands are all calculated with the estimated relative positions. Furthermore, to handle the problem that each robot may encounter new neighboring robots which are not localized, the localization-enhancing command $v_{i}^{\rm enh}$ is designed to enhance the relative localization among adjacent robots (see Fig. 2C). Simulation and experiment results demonstrate that the robots can form desired shape using only local distance and displacement measurements in the absense of external localization infrastructures.

### v4

@@ Abstract
In this article, we address the shape formation problem for massive robot swarms in environments where external localization systems are unavailable. Achieving this task effectively with solely onboard measurements is still scarcely explored and faces some practical challenges. To solve this challenging problem, we propose the following novel results. Firstly, to estimate the relative positions among neighboring robots, a concurrent-learning based estimator is proposed. It relaxes the persistent excitation condition required in the classical ones such as the least-square estimator. Secondly, we introduce a finite-time agreement protocol to determine the shape location. This is achieved by estimating the relative position between each robot and a randomly assigned seed robot. The initial position of the seed one marks the shape location. Thirdly, based on the theoretical results of the relative localization, a novel behavior-based control strategy is devised. This strategy not only enables the adaptive shape formation of large groups of robots but also enhances the observability of inter-robot relative localization. Numerical simulation results are provided to verify the performance of our proposed strategy compared to the state-of-the-art ones. Additionally, outdoor experiments on real robots further demonstrate the practical effectiveness and robustness of our methods.
@@
Shape formation has a broad potential for large groups of robots to execute certain tasks, such as object transport, forest firefighting, and entertainment shows. However, most of the existing approaches rely on external localization infrastructures, rendering them impractical in environments where such systems are not available. To address this issue, this article proposes an integrated strategy that can achieve shape formation for large groups of robots by using local distance and displacement measurements. This strategy consists of three main components. Firstly, a relative localization estimator is introduced to estimate the relative positions among neighboring robots. Secondly, a protocol for reaching a consensus on the desired shape’s position is proposed. Thirdly, a behavior-based controller is developed to achieve massive shape formation and enhance the observability of relative localization. More details of the proposed algorithms and swarm robotic systems are provided in this article.
@@ I Introduction
Shape formation has received significant attention in recent years due to its broad potential across many applications [1, 2]. The objective of shape formation is to guide the robot swarms from an initial configuration to form a user-specified shape through local interactions. To achieve this objective, various control methods have been explored, including the ones based on artificial potential [3], edge following [4], mean-shift exploration [5, 6], and goal assignment [7]. However, most of the methods rely on external localization systems such as GPS, to localize robots, which can be restrictive for practical applications [8]. For instance, external localization systems may be impractical or infeasible in some scenarios, such as indoor environment, underwater, or building atrium. Rather than relying on external localization systems, robots equipped solely with onboard sensors are more autonomous.
@@
In this article, we aim to investigate the problem of relative localization in shape formation using only onboard measurements, as shown in Fig. 1. To handle this problem, the existing results can be roughly classified into the following two categories. The first class is based on the rigidity conditions. These methods usually require an assumption that the measurement topology among robots satisfy rigidity conditions like distance rigidity [9, 10] or bearing rigidity [11, 12]. The second class, which has also attracted research attention, is rigidity-free methods. To that end, researchers incorporate the persistent excitation (PE) characteristics into the control command, such as noise [13, 14] or time-varying reference [15, 8], to relax rigidity conditions on the measurement topology. Although rigidity and PE conditions are necessary in some collaborative tasks, they may not be essential in the context of massive shape formation. More significantly, imposing these conditions may bring a series of technical challenges, as outlined below.
@@
Firstly, the rigidity condition is the foundation of the global convergence property of most existing results [9, 10, 11, 12]. The rigidity condition indicates that any two neighboring robots must involve onboard measurements and the measurement topology for the entire robot swarms should be rigid. However, the two robots that are topological neighbors may be far apart physically, which brings difficulties in measurement and communication between them. Furthermore, the rigidity condition across the entire topology is also difficult to guarantee, especially for large groups of robots.
@@
Secondly, to relax the rigidity condition, researchers introduce the PE condition based localization estimator. In practice, the most common way to fulfill the PE condition is embedding noise or time-varying reference into the control commands of individual robots [13, 14]. However, these PE characteristics may propagate within robot swarms due to interactions between robots. This can result in non-smooth motion trajectories for robot swarms, leading to deteriorated control performance in real-world applications.
@@
Motivated by these challenges, we focus on the shape formation control for large groups of robots without PE condition. The contributions of this article are summarized below.
@@
1) To achieve inter-robot relative localization without the PE condition, a concurrent-learning based relative position estimator is firstly introduced. Compared with the existing results [14, 13, 15], both current and historical measurement data are utilized in the estimator. As a result, it is not necessary to incorporate PE characteristics into the control command, resulting in smoother motion trajectories for the robots. Theoretical analysis results are provided to guarantee the convergence of the estimation error. Furthermore, we provide the optimal strategy for relative motion and data collection between robots. We demonstrate that the proposed estimator achieves the largest convergence rate with the optimal strategy.
@@
2) To achieve the shape localization agreement for the entire swarms in the absence of global coordinates, the location of the shape center is selected as the initial position of a randomly assigned seed robot. Note that the position of the seed robot is unavailable and irrelevant. By transforming the global-coordinate based shape agreement problem [5] into a relative position estimation problem, we propose a finite-time consensus-based agreement protocol. This approach eliminates the need for the rigidity condition on the measurement topology [9, 10, 11, 12] to achieve shape localization agreement. Theoretical analysis results are also provided to guarantee the convergence of the proposed method.
@@
3) For the shape formation control of the robot swarms, we design a novel behavior-based control strategy which is based on the theoretical insights from relative localization. This strategy enables shape formation in large groups of robots in the absence of external localization system. Additionally, it enhances the observability of relative localization among robots. Extensive simulation results are provided to verify the effectiveness of the proposed methods. Moreover, compared with the existing swarm shape formation methods based on relative localization [8, 16, 17], outdoor experiments with a number of robots up to six are also provided to verify the effectiveness, stability, and robustness of the proposed method.
@@
The rest of this article is organized as follows. Section II provides the related works. Section III introduces an overview of the entire system, adaptive relative localization method and behavior-based shape formation control strategy are proposed in Section IV and Section V. Simulation and experimental results are discussed in Section VI and Section VII. Finally, Section VIII concludes this article.
@@ II-A Distributed Shape Formation
Traditional researches on swarm formation control rely on consensus and graph theory [18, 19, 20]. In these methods, each robot achieves shape formation by tracking a predefined desired reference in a distributed manner. However, the neighboring robots defined in the desired formation may be far apart physically, which is unreasonable in real applications.
One class of shape formation strategies that are widely studied is based on goal assignment. In these methods, each robot is dynamically assigned a unique goal location using decentralized approaches such as distributed auction [21] and task-swapping algorithm [7]. Once the robots are assigned unique goal locations, the subsequent task is simply to guide the robots to reach their respective goal locations. Many existing results, such as distributed model predictive control [21], optimal reciprocal collision avoidance [22] and planning based methods [23], can accomplish this task effectively.
The other class of strategies are based on local behavior [24, 25, 26]. Inspired by natural processes like embryogenesis, the method in [27] can spontaneously generate swarm shapes with a reaction-diffusion network. Similarly, a random walk based method is proposed in [1]. These methods exhibit significant randomness in both the shape formed and the robots’ motion trajectories. Another common assignment-free methods for shape formation are based on artificial potential fields (APF). However, APF-based methods often encounter a well-known locally minimum problem. To address this issue, an information integration and feedback based method is proposed in [3]. However, relying solely on APF cannot effectively avoid the deadlock phenomena. In our previous work [5], the idea of mean-shift exploration is adopted in the shape formation of robot swarms, which significantly improves the efficiency of shape formation, especially for large groups of robots.
All these methods are based on the robots’ global coordinates. However, one can see that many simple organisms in nature can achieve complex shapes with only basic measurement information rather than global coordinates. Consequently, the development of shape formation techniques for robot swarms using only local measurements remains significant.
@@ II-B Relative Localization Estimation
Relative localization between robots is a fundamental problem in swarm robot research. Different from traditional anchor-based localization methods [28, 29], we focus on the relative localization strategy which does not rely on external anchors [30]. In recent years, visual-measurement based relative localization schemes have received significant attention. The work in [31] proposes a fisher information matrix (FIM) based localization-enhanced method with bearing measurement. In [32] and [33], partially observations based framework is proposed to estimate the relative poses within a robotic swarm. To handle the limited camera’s field of view, an omnidirectional cameras based simultaneous localization and mapping system named $D^{2}$ is proposed in [34]. However, the vision-based relative localization methods may lead to high computational resource consumption [16].
In contrast, thanks to the availability of low-cost ultra wide band (UWB) ranging sensors, distance based localization schemes are also widely explored. A common approach is to install multiple UWB modules on the robot. In the lastest results [35] and [36], the relationship between the motion trajectory and localization performance of robot is studied. However, the use of multiple UWB sensors may increase the structural complexity of the robot. In [14], a distance and displacement measurement based relative position estimator is proposed, requiring only one UWB sensor. However, this estimator only converges when the relative motion trajectory between robots satisfies the PE condition. As a result, the periodic noise signal is added to the control command, which is similar to [13]. In [16], based on the FIM theory, the authors analyze how the relative motion between two robots can enhance relative localization. In the latest result [37], the PE condition is satisfied through the rotation of a UWB module on each robot. Nevertheless, the design of the turntable adds complexity to the robot’s structure.
@@ II-C Local Measurement Based Shape Formation
To achieve shape formation tasks with only local measurements, researchers in control and robotics have explored various strategies. One common strategy is based on a trilateral localization algorithm. In [4], a distance measurement based shape formation strategy is proposed. In this work, each robot localizes itself by measuring distance with at least four localized robots. However, this method requires the placement of anchor robots and only edge robots can move to form shapes, resulting in lower efficiency. Similarly, multiple ranging sensors are used for inter-robot localization in [38]. In [39], the elastic link is adopted to enhance localization for error-prone individuals. However, these methods rely on external facilities, making practical implementation challenging.
@@
In summary, while the existing approaches offer promising solutions for local measurement based swarm shape formation control, they all have limitations. Currently, there is still no method that achieves efficient shape formation control for large groups of robots without relying on the PE condition. The distributed shape formation problem with solely onboard measurements remains a challenge yet to be fully addressed.
III Preliminaries and System Overview
III-A Notions
TABLE I: List of main symbols used in this article
Symbols Description
$t_{k}$pbox10cmUWB & IO measurement time instant for robot $i$.
$p_{i}(t_{k})$pbox10cmPosition of the robot $i$.
$u_{i}(t_{k})$pbox10cmDisplacement of robot $i$ from time instant $t_{k}$ to $t_{k+1}$.
$z_{i}(t_{k})$pbox10cmDisplacement of robot $i$ from initial time instant $t_{0}$ to $t_{k}$.
$d_{ij}(t_{k})$pbox10cmDistance between robots $i$ and $j$.
$t_{{\rm c},m}$pbox10cmStarting time instant of the $m$-th measurement data collection.
$y_{ij}$pbox10cmAuxiliary signal which records innovation.
$\hat{p}_{ij}$pbox10cmEstimate of real-time relative position between robots $i$ and $j$.
$\hat{q}_{i,0}$pbox10cmEstimate of initial relative position between robot $i$ and the
seed robot 00.
$v_{i}$pbox10cmVelocity control command for robot $i$.
@@ III-A Notions
Consider a group of $n+1$ mobile robots in $\mathbb{R}^{2}$ and $n\geq 2$. Let $p_{i}\in\mathbb{R}^{2}$ be the position of robot $i=0,1,\ldots,n$ in an inertial frame. The dynamic model of each robot is considered as $\dot{p}_{i}=v_{i}$, where $v_{i}$ is the velocity command to be designed. The velocity command $v_{i}$ satisfies $\lVert v_{i}\rVert\leq v_{\max}$, where $v_{\max}$ is the maximum velocity of each robot. The interaction network among the robots is described by an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, which is composed of a vertex set $\mathcal{V}=\{0,1,\ldots,n\}$ and an edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. If $(i,j)\in\mathcal{E}$, the two robots can receive information from each other and robot $j$ is a neighbor of robot $i$. The neighbor set of robot $i$ is denoted as $\mathcal{N}_{i}=\{j\in\mathcal{V}:(i,j)\in\mathcal{E}\}$. The detailed definition of the neighbor set $\mathcal{N}_{i}$ can be found in V-B. For ease of reading, we have listed the main symbols that appear in this article in Table I.
III-B System Overview
@@
The objective of each robot is to achieve shape formation for the entire robot swarms through local distance and displacement measurements. To achieve this, we design an integration strategy including relative position estimation, shape localization agreement, and shape formation control. The overall system architecture is illustrated in Fig. 2.
@@
The first step involves designing a concurrent-learning based estimator to estimate relative positions to neighboring robots (see Fig. 2A and Section IV). This article enables each robot to achieve relative localization to its neighbors solely through inter-robot distance and displacement measurements which can be obtained using some low-cost sensors such as UWB and inertial odometry (IO), respectively. The proposed distributed adaptive estimators for relative position estimation among neighboring robots are based on the concept of concurrent-learning, i.e. the robots leverage historical and present measurement data to alleviate the need for the PE condition.
@@
The second step is to design an agreement protocol to locate the desired shape which is marked by the initial position of the seed robot (see Fig. 2B and Section V-B). The seed robot plays a stubborn role by insisting on propagating the relative displacement from its initial position (i.e., shape position) to its non-seed neighbors. These non-seed robots utilize this propagated relative displacement to interpret the shape position locally through the proposed agreement protocol, simultaneously propagating their interpretations to other non-seed robots. In this way, the interpretations of non-seed robots gradually converge to the seed one so that all the robots reach a consensus on the localization of the desired shape.
@@
The final step is to devise a behavior-based control law to achieve the shape formation task and, meanwhile, enhance the localization (see Fig. 2C and Section V-C). Once all the robots reach an agreement on the localization of the desired shape, each robot can form the desired shape driven by the proposed behavior-based control law (see Section V). Specifically, the behavior-based control law consists of four velocity commands $v_{i}^{\rm ent}$, $v_{i}^{\rm exp}$, $v_{i}^{\rm int}$, and $v_{i}^{\rm enh}$. The four commands representthe shape-entering, shape-exploring, neighbor-interacting, and localization-enhancing behavior motions, respectively.